Brevin went up last night (6/22) to do a quick servicing on the mirror support compressor, the axis video system, and the oculus camera control computer. In short, all issues are fixed, and the telescope is back to normal operations. Please check images from tonight (6/23) to insure that the telescope is functioning properly. Please reply to this message, or email me personally, if you still find issues. Details on each issue are below.
Oculus camera control computer:
There was another periodic brownout on 6/20. Oculus did not automatically reboot as it was set to do in the BIOS. Cycling the power via the networked APC did not reboot the machine. Not clear at all why this error occurred, but pushing the power button once restored the bios setting, and the machine came up as normal.
Axis Video System:
The video display was showing a large horizontal line for several months. Don Walter visited in May to cycle the power and test a camera or two via the old guider TV monitor, each still showing the long streak. Eventually, the cameras stopped responding completely. We suspected the problem was the AC/DC transformer to the cameras themselves, a very common and inexpensive power plug that only required a bit of stripping– and a very easy test. Brevin installed a new one, and the cameras came back as normal.
Mirror Support Compressor:
Images have shown a typical three-point PSF for the past few weeks, indicative of failed support bags. Brevin found that the motor circuit breaker had tripped and reset it. In the dome, they also found that one of the release valves (a ball and nut unit) had fallen off, perhaps loose over some time. They replaced the value, and the bags came back to nominal pressure (10 psi when vertical, < 5 psi at horizons).